# Pre-training Cross-lingual Open Domain Question Answering with Large-scale Synthetic Supervision

### v1

@@ 4 Experiments
4.1 Dataset
We evaluate our approach on the Xor-TyDi QA dataset (Asai et al., 2021a), where the XOR-Retrieve task involves cross-lingual retrieval and QA and XOR-Full centers on multilingual open-retrieval QA. For both tasks, we employ MKQA (Longpre et al., 2021), a dataset generated by translating questions from Natural Questions (NQ) (Kwiatkowski et al., 2019), to evaluate our model’s zero-shot cross-lingual transfer ability on unseen languages. We use the February 2019 English Wikipedia dump as $\bm{C^{En}}$ and use the same Wikipedia dumps of 13 diverse languages from all 7 languages of Xor-TyDi QA and a subset of MKQA languages as $\bm{C^{Multi}}$ (Asai et al., 2021a).
We take R@2kt and R@5kt as evaluation metrics for cross-lingual retrieval, which measures the fraction of queries for which the top-$n$ retrieved tokens contain the answer. The average R@2kt and R@5kt scores across all languages are used for model comparison. F1 and Exact Match (EM) are used for evaluating question-answering tasks, with token-level BLEU scores supplemented on XOR-Full and MKQA. The average F1 score across all languages is used for performance comparisons.
@@ 4.2 Experimental Settings
Pre-training Corpus
In stage-1, for each $\mathcal{W}^{En}\in\bm{C^{En}}$, we gather its parallel pages across various languages. We consider 15 distinct languages, comprising 7 from Xor-TyDi QA, and 8 are high-resource or closely related to the 7 evaluated languages. Parallel sentences are mined from each pair of parallel pages. A state-of-the-art NER tagger is applied to each English sentence, and we retain pairs featuring contained named entities.
In stage-2, data generation is limited to 7 languages on Xor-TyDi QA. We employ LLaMA-2-7B to generate one transformed question per training example with 3 randomly sampled meta-examples in the same language as the prompt. We generate multiple questions for each example in low-resource languages. More details are in Appendices A.1.1 and A.1.2.
Рис. 2: The training pipeline of our method.
Training Recipe
An English teacher $\theta^{En}$ is first trained on NQ (Kwiatkowski et al., 2019) as in Jiang et al. (2022). Undertaking stage-1 pre-training by enforcing mt5-large (Xue et al., 2021) to learn from $\theta^{En}$ results in CLASS-US-Stage1, followed by stage-2 pre-training to obtain the unsupervised model, CLASS-US. The zero-shot model CLASS-ZS is acquired by training CLASS-US on English data (i.e., NQ), followed by fine-tuning on labelled data from Xor-TyDi QA to obtain the supervised model CLASS. An overview is shown in Figure 2 and more details are in Appendix A.1.3.
Method Size R@2kt R@5kt
@@
CLASS-US 410M 66.0 75.7 63.4 57.7 63.5 68.8 70.6 66.5 71.2 81.6 69.4 66.8 70.5 75.1 77.3 73.1
Zero-shot Retrievers
${}^{\dagger}$ 220M 43.4 53.9 55.1 40.2 50.5 30.8 20.2 42.0 52.4 62.8 61.8 48.1 58.6 37.8 32.4 50.6
@@
ReAtt+MT 583M 63.1 67.7 20.7 55.9 60.3 55.3 58.4 54.5 67.3 71.0 29.3 61.8 67.0 61.2 66.4 60.6
CLASS-ZS 410M 65.1 79.3 67.8 60.6 61.1 69.2 74.4 68.2 72.5 83.2 73.9 70.5 69.1 75.1 81.9 75.2
(Semi-) Supervised Retrievers
CORA 557M 32.0 42.8 39.5 24.9 33.3 31.2 30.7 33.5 42.7 52.0 49.0 32.8 43.5 39.2 41.6 43.0
${}^{\dagger}$ 557M 38.8 48.4 52.5 26.6 44.2 33.3 39.9 40.5 48.9 60.2 59.2 34.9 49.8 43.0 55.5 50.2
[…]
QuiCK 557M 52.8 70.1 62.2 54.8 62.8 57.8 70.6 61.3 63.8 78.0 65.3 63.5 69.8 67.1 74.8 68.9
${}^{\ast}$ 278M - - - - - - - 66.0 70.2 85.9 69.4 65.1 68.8 68.8 83.2 73.1
[…]
CLASS 410M 67.3 80.9 67.2 64.7 71.6 69.6 79.8 71.6 74.8 84.5 72.3 73.9 79.3 77.2 85.3 78.2
Method Size F1 EM
Ar Bn Fi Ja Ko Ru Te Avg Ar Bn Fi Ja Ko Ru Te Avg
${}^{\dagger}$ 835M 17.9 19.4 24.5 13.1 14.3 17.2 14.4 17.2 11.7 14.1 18.3 10.7 9.9 11.1 10.2 12.3
${}^{\dagger}$ 28.0 25.6 29.3 19.2 19.4 18.4 3.8 20.5 23.4 20.3 22.1 13.8 14.2 13.6 2.7 15.7
ReAtt+MT 1.19B 26.0 36.6 3.5 18.4 29.3 29.3 27.0 24.3 17.1 30.1 1.7 14.5 22.7 22.8 21.0 18.6
${}^{\dagger}$ 39.5 42.1 28.2 23.5 30.5 34.8 31.6 32.9 28.5 34.4 21.3 17.4 23.8 26.4 25.1 25.3
CLASS 1.23B 44.0 59.4 43.9 40.6 44.8 47.9 50.1 47.2 35.7 52.2 35.3 35.1 37.1 38.8 44.1 39.8
[…]
Method Size F1 Macro Average
Ar Bn Fi Ja Ko Ru Te F1 EM BLEU
${}^{\dagger}$ 31.1 21.9 21.4 12.4 12.1 17.7
${}^{\dagger}$ 7.2 4.3 17.0 7.9 7.1 13.6 0.5 8.2 3.8 6.8
ReAtt+MT 1.19B 15.0 10.5 1.8 13.1 14.9 15.4 8.2 11.3 5.5 9.5
${}^{\dagger}$ 18.0 29.1 13.8 5.7 15.2 14.9 15.6 16.0 9.9 14.9
${}^{\dagger}$ 15.8 9.6 20.5 12.2 11.4 16.0 0.5 17.3 7.5 10.7
${}^{\dagger}$ 1.14B 42.9 26.9 41.4 36.8 30.4 33.8 30.9 34.7 25.8 23.3
CLASS 1.23B 49.5 32.0 49.6 44.7 37.5 41.4 42.0 42.4 32.7 29.2
Incomparable Models
[…]
4.3 Main Results
XOR-Retrieve
Table 1 shows the results on the dev set of XOR-Retrieve. CLASS, which exclusively employs question-answer pairs for training, demonstrates a substantial performance advantage over all baselines that rely on passage labels for contrastive learning. This advantage is particularly pronounced under unsupervised and zero-shot settings, where both variants, CLASS-US and CLASS-ZS, achieve improvements of more than 10% compared to state-of-the-art methods. In the English span extraction task, CLASS surpasses all competitive baselines by a significant margin (+15%), which is attributed to its superior retrieval abilities in gathering more relevant documents.
XOR-Full
Table 2 reports the results of CLASS on XOR-Full. CLASS achieves superior performance when compared to a series of baseline models and the prior state-of-the-art CORA model in all tested languages, showcasing an average improvement of 7.8%. Compared to methods that rely on machine translation to generate a substantially larger pool of multilingual training data from English datasets, CLASS is comparable to Sentri but falls behind LAPCA. [^5] A direct comparison with Sentri and LAPCA is not feasible since the Wikipedia pages they employed as knowledge sources are different from ours and Asai et al. (2021b). The most pronounced performance gaps are in Bengali and Korean, with the fewest two training samples available within XOR-Full. We believe it is the translated questions used by Sentri and LAPCA that alleviate such discrepancies. We expect CLASS can be further improved by introducing such augmented training data during fine-tuning, which we leave as future work.
Рис. 3: Zero-shot cross-lingual retrieval results (Avg. R@2kt) on unseen languages of MKQA.
MKQA
We assess the zero-shot performance of CLASS in various unseen languages included in MKQA. Figure 3 shows that in cross-lingual retrieval tasks, all variants of our method exhibit promising results. Notably, CLASS-US can outperform the supervised model CORA significantly, and further fine-tuning on English data leads to substantial improvements. Interestingly, CLASS underperforms CLASS-ZS, despite being further fine-tuned on multilingual data. We attribute this phenomenon to two factors: the limited number of queries in XOR-Retrieve may result in the model overfitting to these specific languages; the query topics differ, as MKQA is the literal translation of NQ while XOR-Retrieve is created independently. In the multilingual QA task, we observe similar patterns where CLASS-ZS achieves the best zero-shot performance across unseen languages while further fine-tuning this model on XOR-Full data results in a reduction in its overall generalizability. [^6] We leave the exploration of improved fine-tuning techniques that strike a balance between enhancing supervised and zero-shot cross-lingual results as future work. Detailed results in each language are in Appendix B.
4.4 Analysis
We present detailed ablation studies in this section. More analyses are in Appendix C.
Рис. 4: Ablations on stage-1 pre-training, with results on the XOR-Retrieve dev set reported. ${}^{\ast}$ indicates unseen languages from MKQA.
Ablations
We compare CLASS with different variants to study the impact of different components in stage-1 pre-training. As shown in Figure 4, we found that: i) Removing queries either in English (-$q^{En}$) or in target languages (-$q^{L}$) leads to performance degradation. ii) The cross-lingual alignment regularisation (-$\mathcal{L}^{align}$) enhances the performance of the model trained with parallel queries by ensuring consistent predictions. iii) Pre-training solely on the 7 languages of Xor-TyDi QA (w/ 7 langs) does not significantly impact average performance but affects specific low-resource languages. In particular, adding data from languages related to Telugu and Japanese (e.g., Tamil & Chinese) yields improvements. Moreover, including a wider range of languages improves generalisation to unseen languages. For instance, pre-training in German enhances understanding of the West Germanic languages (i.e., Danish, Dutch, and Norwegian). iv) When comparing the approaches of gathering parallel queries, our method outperforms code-switching (w/ CS), which creates pseudo-translations through lexicon replacement based on bilingual dictionaries, and machine translations (w/ MT). This inferiority is primarily attributed to the limited coverage of bilingual dictionaries and poor translation quality in low-resource languages.
Method XOR-Retrieve XOR-Full
R@2kt R@5kt F1 EM F1 EM BLEU
Unsupervised
CLASS-US (AB) 66.5 73.1 21.7 17.0 18.4 12.0 14.6
  - Stage-2 (A) 59.1 67.4 6.0 4.1 5.7 3.9 4.0
  - Query TF (AC) 66.1 73.1 7.4 5.5 7.2 4.8 4.9
Zero-shot
CLASS-ZS (ABD) 68.2 75.2 34.5 28.0 23.9 15.8 19.4
  - Stage-2 (AD) 62.9 71.1 32.8 26.9 13.7 8.1 8.3
  - Pre-train (D) 27.6 36.3 17.0 13.0 15.4 9.6 11.0
Supervised
CLASS (ABDE) 71.6 78.2 47.2 39.8 42.4 32.7 29.2
  - Stage-2 (ADE) 69.6 75.7 46.0 38.7 42.5 33.1 29.1
  - Pre-train (DE) 62.8 69.3 41.6 35.2 41.9 32.6 28.7
Таблица 3: Effects of two-stage pre-training. Results on the dev sets are reported. Symbols within brackets are described in Figure 2.
Effects of Two-stage Pre-training
We evaluate the effectiveness of our proposed pre-training framework by excluding either the stage-2 or the entire pre-training procedures. Table 3 showcases the performance on both XOR-Retrieve and XOR-Full across unsupervised, zero-shot, and supervised settings. The inclusion of stage-2 pre-training dramatically boosts the performance in both unsupervised and zero-shot scenarios. Removing Query Transformation in stage-2 by solely employing cloze-style questions has no discernible impact on retrieval performance but yields sub-optimal QA results. This highlights the importance of generating queries that resemble the formats of natural questions to downstream QA tasks. When discarding the entire pre-training process, wherein the model only uses the initial retrievals from DPR (Karpukhin et al., 2020) for warm-up training, we observe a notable decline in performance on both datasets. In supervised settings, the advantages conferred by our proposed pre-training diminish in the presence of labelled data. These benefits become progressively inconsequential as the number of supervised data increases (i.e., 15K in XOR-Retrieve v.s. 61K in XOR-Full).

### v3

@@ 4 Experiments
Datasets, Baselines and Metrics.
We evaluate our model on the Xor-TyDi QA dataset (Asai et al., 2021a), with XOR-Retrieve for cross-lingual retrieval, and XOR-Full for multilingual open-domain QA. We employ MKQA (Longpre et al., 2021) for zero-shot evaluation on unseen languages. We use the February 2019 English Wikipedia dump as $\bm{C^{\textbf{En}}}$ and use the Wikipedia dumps of the same date, consisting of 13 diverse languages from all 7 languages of Xor-TyDi QA and a subset of MKQA languages as $\bm{C^{\textbf{Multi}}}$ (Asai et al., 2021a).
We compare retrieval performance with translate-test methods DPR+MT (Asai et al., 2021a), multilingual dense passage retrievers mDPR, CORA, Sentri, QuiCK, LAPCA, SWIM-X (Asai et al., 2021a, b; Sorokin et al., 2022; Ren et al., 2022; Abulkhanov et al., 2023; Thakur et al., 2023), and multi-vector retriever DrDecr (Li et al., 2022). We report top-$n$ retrieval accuracy, the fraction of queries for which the top-$n$ retrieved tokens contain the answer. We compare QA results with multilingual models that use BM25 for monolingual retrieval, translate-test models MT+DPR, GMT+GS, MT+Mono and ReAtt+MT (Asai et al., 2021a; Jiang et al., 2022), and multilingual fusion-in-decoder models CORA, Sentri and LAPCA using F1, exact match (EM) and BLEU scores.
@@ 4.1 Experimental Settings
Pre-training Corpus.
In cross-lingual retrieval pre-training, we gather the parallel pages across various languages for each $\mathcal{W}^{En}\in\bm{C^{\textbf{En}}}$. We consider 15 distinct languages, with 7 from Xor-TyDi QA and 8 being high-resource or closely related to the 7 evaluated languages. Parallel sentences are mined from each pair of parallel pages. A state-of-the-art NER tagger is applied to each English sentence, and we retain pairs that contain named entities.
In multilingual QA pre-training, data generation is limited to 7 languages on Xor-TyDi QA. We employ LLaMA-2-7B to generate one transformed question per training example with 3 randomly sampled meta-examples in the same language as the prompt. We generate multiple questions for each example in low-resource languages. More details are in Appendices A.1.1 and A.1.2.
Рис. 3: Our training pipeline. CLR: cross-lingual retrieval, MLQA: multilingual question answering, QT: query transformation, PT: pre-training, FT: fine-tuning.
Method # Total Params Pre-train Data Fine-tuning Data R@2kt R@5kt
@@
LAPCA§ 560M Wikipedia 51.1 50.2 48.6 35.1 57.3 32.2 64.4 48.4 61.0 58.4 52.6 40.5 66.7 40.8 70.1 55.7
SWIM-X 580M mC4 SWIM-IR 50.8 65.1 56.1 48.1 54.0 55.7 66.4 56.6 57.9 75.0 65.6 59.3 58.9 64.6 74.4 65.1
CLASS-US 410M Wikipedia 66.0 75.7 63.4 57.7 63.5 68.8 70.6 66.5 71.2 81.6 69.4 66.8 70.5 75.1 77.3 73.1
w/ Dense 410M Wikipedia 54.4 67.4 58.6 47.7 51.6 59.9 65.6 57.9 64.8 73.0 64.7 57.3 58.6 67.9 70.6 65.3
Zero-shot Retrievers
DPR+MT 220M NQ 43.4 53.9 55.1 40.2 50.5 30.8 20.2 42.0 52.4 62.8 61.8 48.1 58.6 37.8 32.4 50.6
LAPCA§ 560M Wikipedia NQ+XPAQ 46.2 50.3 56.6 41.4 48.7 52.3 54.6 50.0 53.0 60.5 66.2 49.7 56.1 60.7 63.8 58.6
ReAtt+MT 583M NQ 63.1 67.7 20.7 55.9 60.3 55.3 58.4 54.5 67.3 71.0 29.3 61.8 67.0 61.2 66.4 60.6
CLASS-ZS 410M Wikipedia NQ 65.1 79.3 67.8 60.6 61.1 69.2 74.4 68.2 72.5 83.2 73.9 70.5 69.1 75.1 81.9 75.2
w/ Dense 410M Wikipedia NQ 59.2 70.1 59.9 51.5 57.2 51.5 72.3 60.2 66.7 78.6 66.6 60.2 63.2 58.2 78.2 67.4
(Semi-) Supervised Retrievers
CORA 557M NQ+XOR 32.0 42.8 39.5 24.9 33.3 31.2 30.7 33.5 42.7 52.0 49.0 32.8 43.5 39.2 41.6 43.0
mDPR 557M NQ+XOR 38.8 48.4 52.5 26.6 44.2 33.3 39.9 40.5 48.9 60.2 59.2 34.9 49.8 43.0 55.5 50.2
Sentri§ 560M NQ+TQA+XOR 47.6 48.1 53.1 46.6 49.6 44.3 67.9 51.0 56.8 62.2 65.5 53.2 55.5 52.3 80.3 60.8
QuiCK 557M NQ+XOR 52.8 70.1 62.2 54.8 62.8 57.8 70.6 61.3 63.8 78.0 65.3 63.5 69.8 67.1 74.8 68.9
DrDecr 278M WikiMatrix NQ+XOR - - - - - - - 66.0 70.2 85.9 69.4 65.1 68.8 68.8 83.2 73.1
LAPCA§ 560M Wikipedia NQ+XPAQ+XOR 61.1 76.9 72.6 60.9 69.1 69.1 75.6 69.3 70.2 83.8 79.6 69.7 73.6 75.5 83.1 76.5
CLASS 410M Wikipedia NQ+XOR 67.3 80.9 67.2 64.7 71.6 69.6 79.8 71.6 74.8 84.5 72.3 73.9 79.3 77.2 85.3 78.2
w/ Dense 410M Wikipedia NQ+XOR 66.7 79.6 64.3 58.1 66.0 64.1 77.7 68.1 70.6 84.9 71.0 66.0 72.6 70.0 81.9 73.9
Таблица 1: Results on the dev set of XOR-Retrieve. The best and second-best results are marked in bold and underlined. $\dagger$ denotes results reported by Asai et al. (2021a). $\ast$ indicates human-translated supervised parallel queries released by XOR-Retrieve are used for training. §$\S$ represents methods that employ MT systems for training data augmentation.
Method # Total Params Pre-training Data Fine-tuning Data F1 Macro Average
Ar Bn Fi Ja Ko Ru Te F1 EM BLEU
BM25 XOR 31.1 21.9 21.4 12.4 12.1 17.7
MT+DPR NQ 7.2 4.3 17.0 7.9 7.1 13.6 0.5 8.2 3.8 6.8
ReAtt+MT 1.19B NQ 15.0 10.5 1.8 13.1 14.9 15.4 8.2 11.3 5.5 9.5
GMT+GS NQ 18.0 29.1 13.8 5.7 15.2 14.9 15.6 16.0 9.9 14.9
MT+Mono NQ+XOR 15.8 9.6 20.5 12.2 11.4 16.0 0.5 17.3 7.5 10.7
CORA 1.14B NQ+XOR 42.9 26.9 41.4 36.8 30.4 33.8 30.9 34.7 25.8 23.3
CLASS 1.23B Wikipedia NQ+XOR 49.5 32.0 49.6 44.7 37.5 41.4 42.0 42.4 32.7 29.2
    w/ Dense 1.23B Wikipedia NQ+XOR 49.1 32.0 46.7 44.1 38.4 39.9 41.1 41.6 32.5 28.2
Incomparable Models (for Reference)
Sentri§ 1.14B NQ+TQA+XOR 52.5 31.2 45.5 44.9 43.1 41.2 30.7 41.3 34.9 30.7
LAPCA§ 1.14B Wikipedia NQ+XPAQ+XOR 53.4 50.2 49.3 44.7 49.5 49.3 38.9 47.8 38.7 35.5
Таблица 2: QA results on the XOR-Full dev set. The best and second-best results are marked in bold and underlined. $\dagger$ denotes results from  Asai et al. (2021b). §$\S$ indicates methods that use synthetic and translated English datasets.
Training Sequence.
Figure 3 shows the complete pre-training and fine-tuning sequence. i) Cross-lingual Retrieval Pre-training (CLR-PT): We pre-train mt5-large (Xue et al., 2021) as in §3.1 to get CLASS-US-Stage1, with English teacher being ReAtt (Jiang et al., 2022) trained on NQ (Kwiatkowski et al., 2019). ii) Multilingual QA Pre-training (MLQA-PT): CLASS-US-Stage1 is further pre-trained as in §3.2 to obtain the unsupervised CLASS-US. iii) Fine-tuning: We first fine-tune CLASS-US on NQ to obtain the zero-shot CLASS-ZS, which is then trained on supervised data from Xor-TyDi QA to derive CLASS. We use the same training objective $\mathcal{L}_{\text{e2e}}$ as in MLQA-PT.
4.2 Main Results
XOR-Retrieve.
Table 1 shows the results on the dev set of XOR-Retrieve. CLASS, which exclusively employs question-answer pairs for training, demonstrates a substantial performance advantage over all baselines that rely on passage labels for contrastive learning. This advantage is particularly pronounced under unsupervised and zero-shot settings, where both variants, CLASS-US and CLASS-ZS, achieve improvements of more than 10% over state-of-the-art methods ($p<0.001$). [^4] Paired Student’s t-test (Dror et al., 2018). The Dense Retrieval variant (i.e., w/ Dense) consistently outperforms other competitive baselines and is comparable to LAPCA with only 73% of the parameters. This highlights that our approach is versatile and can be applied to enhance various kinds of retrievers.
XOR-Full.
Table 2 reports the results of CLASS on XOR-Full. Both CLASS and the variant employing dense retrieval achieve superior performance when compared to a series of baseline models and the prior state-of-the-art CORA model in all tested languages, showcasing an average improvement of up to 7.8% ($p<0.001$). Compared to methods that rely on machine translation to generate a substantially larger pool of multilingual training data from English datasets, CLASS is comparable to Sentri but falls behind LAPCA. [^5] A direct comparison with Sentri and LAPCA is not feasible since the Wikipedia pages they employed as knowledge sources are different from ours and Asai et al. (2021b). The most pronounced performance gaps are in Bengali and Korean, with the fewest two training samples available within XOR-Full. We believe it is the translated QA pairs used by Sentri and LAPCA that alleviate such discrepancies, and further improvements are expected when integrating such augmented data.
Рис. 4: Zero-shot cross-lingual retrieval and multilingual QA results on unseen languages of MKQA. Macro average results across all test languages are reported. Languages included are: Da, De, Es, Fr, He, Hu, It, Km, Ms, Nl, No, Pl, Pt, Sv, Th, Tr, Vi, Zh-cn, Zh-hk, and Zh-tw.
MKQA.
We assess the zero-shot performance of CLASS in various unseen languages included in MKQA. Figure 4 shows that in cross-lingual retrieval tasks, all variants of our method exhibit promising results. Notably, CLASS-US surpasses the supervised model CORA significantly, and further fine-tuning on English data leads to substantial improvements. Interestingly, CLASS underperforms CLASS-ZS, despite being further fine-tuned on multilingual data. Similar patterns are observed in the multilingual QA task, where CLASS-ZS achieves the best zero-shot performance across unseen languages while supervised fine-tuning on XOR-Full hurts the generalisability. We attribute this phenomenon to three factors: 1) the limited number of queries in Xor-TyDi QA leads to overfitting to these specific languages, as we observe that the model’s performance in both retrieval and QA tasks of MKQA decreases as the fine-tuning on Xor-TyDi QA continues; 2) the query topics differ, as MKQA was translated from NQ while Xor-TyDi QA questions were created by native speakers in target languages; 3) the answer type differs (free spans on Xor-TyDi QA v.s. WikiData aligned entities on MKQA). Our manual inspection reveals that CLASS is more likely to generate free-span answers than CLASS-ZS. Detailed results in each language are in Appendix B Tables 5 and 6.
Size Ar Bn Fi Ja Ko Ru Te Avg.
CLASS-ZS 1.23B 26.8 22.9 20.3 23.1 27.2 25.0 21.9 23.9
Gemma 7B 13.4 19.0 21.7 20.2 20.5 23.0 23.4 20.2
LLaMA3 8B 22.7 13.2 22.9 17.8 19.0 19.2 28.9 20.5
CLASS 1.23B 32.3 28.1 29.9 25.7 29.5 27.7 24.7 29.8
Таблица 3: F1 scores on XOR-Full under 5-shot learning settings. Gemma and LLaMA3 are RALM baselines. CLASS is obtained through 5-shot fine-tuning over CLASS-ZS.
4.3 Few-shot Results
To further demonstrate the superiority of CLASS under low-resource settings, we compare our method against retrieval-augmented language model (RALM) baselines, where all systems are provided 5-shot supervision. The shots are sampled from the XOR-Full training data. For the RALM baselines we prompt a LLM (Llama3-8B (Dubey et al., 2024) and Gemma-7B (Team et al., 2024)) with the 5-shot instances and the retrieved passages to the query (retrieved by CLASS-ZS). For CLASS, we fine-tune CLASS-ZS on the same five-shot examples.
As shown in Table 3, CLASS is significantly better than two RALM baselines (+9.3%) despite having roughly 5$\times$ fewer parameters. Meanwhile, note that our zero-shot model CLASS-ZS surpasses the two RALMs by 3.4% (23.9% v.s. 20.5%) without using any supervised data. This demonstrates the superiority of CLASS for zero and low-resource multilingual open-domain QA.
4.4 Analysis
We include quantitative and qualitative error analysis in Appendix C and additional numeric results in Appendix D (Figures 8910)).
Рис. 5: Ablations on cross-lingual retrieval pre-training, with results on the XOR-Retrieve dev set reported. indicates unseen languages from MKQA.
Cross-lingual Retrieval Pre-training Ablations.
We conduct ablation studies to understand the impact of different components in cross-lingual retrieval pre-training, with results shown in Figure 5.
Effects of Learning from Parallel Queries.  Removing queries either in English (-$q^{En}$) or in target languages (-$q^{L}$) leads to performance degradation. Meanwhile, the cross-lingual alignment regularisation (-$\mathcal{L}^{align}$) benefits the model by ensuring consistent predictions across languages.
Comparison with Different Parallel Query Sources.  When comparing the approaches of gathering parallel queries, our method outperforms code-switching (w/ CS), which creates pseudo-translations through lexicon replacement based on bilingual dictionaries, and machine translations (w/ MT). This inferiority is primarily attributed to the limited coverage of bilingual dictionaries and poor translation quality in low-resource languages.
Sensitivity to Pre-training Language.  Removing the extra 8 high-resource languages (w/ 7 langs) does not impact average performance but affects specific low-resource languages in Xor-TyDi QA. In particular, adding languages related to Telugu and Japanese (e.g., Tamil & Chinese) yields improvements. Moreover, including a wider range of languages improves generalisation to unseen low-resource languages with limited parallel Wikipedia links (e.g., adding German data enhances understanding of the West Germanic languages: Danish, Dutch, and Norwegian).
Method XOR-Retrieve XOR-Full
@@ 4.4 Analysis
Unsupervised
CLASS-US (AB) 66.5 73.1 18.4 12.0 14.6 60.0 69.1
- MLQA-PT (A) 59.1 67.4 5.7 3.9 4.0 55.7 74.7
- Query TF (AC) 66.1 73.1 7.2 4.8 4.9 60.1 65.4
Zero-shot
CLASS-ZS (ABD) 68.2 75.2 23.9 15.8 19.4 59.2 69.1
- MLQA-PT (AD) 62.9 71.1 13.7 8.1 8.3 57.0 76.2
- Pre-train (D) 27.6 36.3 15.4 9.6 11.0 52.5 58.6
Supervised
CLASS (ABDE) 71.6 78.2 42.4 32.7 29.2 62.8 78.4
- MLQA-PT (ADE) 69.6 75.7 42.5 33.1 29.1 63.1 77.8
- Pre-train (DE) 62.8 69.3 41.9 32.6 28.7 62.4 71.7
@@
Effects of Two-stage Pre-training.
We evaluate the efficacy of our two-stage proposed pre-training framework. Table 4 showcases the performance on both XOR-Retrieve and XOR-Full under unsupervised, zero-shot, and supervised settings. Integrating multilingual QA pre-training dramatically boosts performance in both unsupervised and zero-shot scenarios. Merely employing cloze-style questions instead of transformed natural questions has minimal impacts on retrieval but yields sub-optimal QA results, highlighting the importance of synthetic natural questions in QA tasks. When discarding the entire pre-training process, we observe a notable drop in both datasets. In supervised settings, the advantages of pre-training diminish with labelled data. This is especially evident in XOR-Full, where the differences between CLASS and the other two variants in QA and in-language retrieval (RL@N) results diminish. While pre-training significantly improves cross-lingual evidence retrieval (RM@N 71.7% -> 78.4%), CLASS does not benefit from this, suggesting its heavy reliance on in-language evidence and inability to reason over cross-lingual evidence when generating answers. See Appendix C for more detailed error analysis.